Circular No. 7955

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002em IN UGC 3430 M. Armstrong, Rolvenden, England, reports his discovery of an appar-

ent supernova (mag 18.0) on an unfiltered CCD image (limiting mag 19.0) taken with a 0.35-m reflector on Aug. 16.07 UT in the course of the U.K. Nova/Supernova patrol. The new object was confirmed at mag 18.0 on an image taken on Aug. 16.95, but it was not present on Armstrong’s images from 2001 Mar. 13 (limiting mag 19.0), 2002 Mar. 7 (limiting mag 19.5), and Apr. 6 (limiting mag 19.0), or on Palomar Sky Survey images taken in 1993 (limiting blue mag 22.5) and 1996 (limiting red mag 20.8). SN 2002em is located at α = 6^{h}15^{m}52^{s}.30, δ = +64^{o}25′49′′.1 (equinox 2000.0), or ≈ 17′′ east and 1′′ south of the center of UGC 3430.

V445 PUPPIS P. A. Woudt, University of Cape Town, reports the following simul-

taneous J, H, and K_{s} photometry of V445 Pup, obtained at the South African Astronomical Observatory, Sutherland, with the 1.4-m IRSF telescope (+ SIRIUS): Mar. 22.763 UT, H = 17.63 ± 0.06, K_{s} = 14.07 ± 0.02; June 19.689, H = 18.09 ± 0.12, K_{s} = 15.29 ± 0.03. V445 Pup was below the detection limit (mag 19.5) in J at both epochs. Woudt adds: “The H–K_{s} color has changed significantly over this period, from +3.56 ± 0.06 to +2.80 ± 0.12, and V445 Pup has become fainter in both photometric passbands. One and a half years after outburst, V445 Pup is still shrouded in a dense dust shell (see IAUC 7730).”

COMET C/2002 O6 (SWAN) Visual m_{1} and coma-diameter estimates: Aug. 4.76 UT, 6.2, 8′ (Y.

Nagai, Yamanashi, Japan, 12×50 binoculars); 6.71, 6.2, 15′ (S. T. Rae, Port Ohope, New Zealand, 8×21 binoculars); 8.12, 6.3, 8′ (S. Baroni, Milan, Italy, 20×80 binoculars); 11.07, 6.0, 14′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 13.11, 6.1, 11′ (M. V. Zanotta, Laino, Italy, 7×42 binoculars); 15.11, 5.9, 11′ (Zanotta).

SUPERNOVA 2002el IN NGC 6986 Corrigendum. On IAUC 7954, line 5, for Oct. 14, read Aug. 14,

2002 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green